OpenStudy (anonymous):

What is the sum of a 22-term arithmetic sequence where the first term is 54 and the last term is -30

OpenStudy (anonymous):

a. 264 b. 288 c. 312 d. 336

OpenStudy (anonymous):

a=54 ,L=-30 Sn=n/2[a+l] s22=22/2[54+-30]=264----->(a)

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

yw!